A chemist burned 150.0 g of C3H8 in the presence of 300.0 g of O2 and obtained 210.43 g of CO2. What is the percent yields of the combustion reaction?

1
Expert's answer
2022-03-13T15:16:43-0400

Solution.

"m(C_3H_8)=150.0g;"

"m(O_2)=300.0g;"

"m(CO_2)=210.43g;"

150.0g 300.0g 210.43g

"C_3H_8+5O_2\to 3CO_2+4H_2O;"

"\nu(C_3H_8)=\dfrac{150.0}{44}=3.41mol;"

"\nu(O_2)=\dfrac{300.0}{32}=9.375mol;"

3.41mol xmol

"C_3H_8+5O_2\to 3CO_2+4H_2O;"

1 5

"x=\dfrac{3.41 mol\sdot5}{1}=17.05;"

9.375mol y

"C_3H_8+5O_2\to 3CO_2+4H_2O;"

5 3

"\nu(CO_2)=\dfrac{9.375\sdot3}{5}=5.625mol;"

"m_t(CO_2)=5.625\sdot44=247.5g;"

"\eta=\dfrac{m(CO_2)}{m_t(CO_2)};"

"\eta=\dfrac{210.43}{247.5}=0.85=85\%;"

Answer: "\eta=85\%."